OpenStudy (anonymous):

if alpha beta and gamma are roots of the equation(x-2)(x^2+6x-11)=0 find (alpha+beta+gamma)

OpenStudy (anonymous):

answer plz

OpenStudy (anonymous):

\[\alpha=2\] \[\beta=\frac{ -6+\sqrt{6^2-4(1)(-11)} }{ 2 }\] \[\Gamma =\frac{ -6-\sqrt{6^2-4(1)(-11)} }{ 2 }\] \[\alpha+\beta+\Gamma=-4\]

OpenStudy (anonymous):

how is alpha 2

OpenStudy (anonymous):

Just to mention, @Jonask shows you a cute property of quadratic equation a^2 + bx + c =0; The sum of the two roots is -b.

OpenStudy (anonymous):

For alpha, x-2 = 0 x = 2 or alpha = 2

OpenStudy (anonymous):

for beta (-b+-root of right??

OpenStudy (anonymous):

For beta and gamma, Jonask wrote that already. But you can find beta + gamma directly, it's equal to -b = -6